# ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning

### v2

@@ 4 Experiments
Pre-training datasets. The datasets are crucial for pre-training a general time series forecasting model. In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc. The details of these datasets are shown in the Appendix A.1.1. To enhance data utilization, we downsample fine-grained datasets to coarser granularity, resulting in approximately 887 million time points.
Evaluation datasets. To conduct comprehensive and fair comparisons for different models, we conducted experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
Baselines. We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$-LLM (Pan et al., 2024). In addition, we select five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b), and Moment (Goswami et al., 2024).
Setup. Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. For fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$. More implementation details are presented in the Appendix A.1.3.
@@ 4.1 In distribution forecasting
Setting. In full-shot setting, we fine-tune pre-trained ROSE and the baselines with the full downstream data. In few-shot setting, we fine-tune all models with only 10% train data.
@@
Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$-LLM
@@
Few-shot results. The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2. ROSE outperforms advanced models when training data is scarce in the target domain.
@@
Figure 4 shows the performance of pre-trained ROSE and ROSE trained from scratch on ETTh1 and ETTm2 with different fine-tuning data percentages, noting the best baselines in full-shot setting. The pre-trained ROSE shows stable, superior performance even with limited fine-tuning samples. Specifically, the pre-trained ROSE exceeds SOTA performance with only 1% train data for ETTh1 and 2% for ETTm2. Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
@@ 4.2 Zero-shot forecasting
Results. As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in Mean Squared Error (MSE). In comparison to Timer and Moirai, ROSE achieves average MSE reductions of 9% and 6%, respectively, and demonstrates a remarkable 43% relative improvement over Moment. Notably, ROSE stands out not only for its superior performance but also for its exceptional lightweight and efficient design, which sets it apart from other foundational models. Detailed analysis of these aspects will be presented in Section 4.3.
Table 2: The results for ROSE and other foundation models in zero-shot setting. The average results of all predicted lengths are listed here. We use ’-’ to indicate that the dataset has been involved in the model’s pre-training, and thus not used for testing.
@@
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
ETTh2 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
ETTm1 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
ETTm2 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
Weather 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
Electricity 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
Traffic 0.588 0.412 0.613 0.407 - - 0.371 0.370 - - 1.411 0.804
@@ 4.3 Model Analysis
As shown in Figure 5, ROSE is a lightweight general model with 7.4M parameters and short inference time, which are only about one-tenth of the second fastest/smallest foundation model (Timer). Importantly, ROSE uses the least number of parameters among foundation models, with its parameter count approaching that of specific models, while exhibiting superior zero-shot performance. This is attributed to our proposed decomposed frequency learning that enhances the model’s comprehension of time series. Concurrently, the TS-Register achieves the adaptive transfer thus efficiently adapting to downstream tasks without the need of scaling up to achieve strong generalizability. Compared to foundation models with large scale, ROSE may better meet the need for general models in real scenarios that require high computational and parameter efficiency as well as high prediction accuracy with scarce downstream data.
@@ 4.4 Ablation studies
Replace Multi-Frequency Masking Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
@@ 5 Conclusion and future work
In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance. ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios. Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities. Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification. We provide our code at https://anonymous.4open.science/r/ROSE-A235.
6 Reproducibility
Our work meets reproducibility requirements. Specifically, you can download our evaluation datasets in a standardized format from the public link (Wang et al., 2024a): https://drive.google.com/drive/folders/13Cg1KYOlzM5C7K8gK8NfC-F3EYxkM3D2, and we provide our code at https://anonymous.4open.science/r/ROSE-A235.
@@ A.1.1 Pre-training Datasets
We use multi-source datasets in pre-training which contain subsets of Monash [Godahewa et al., 2021], UEA [Bagnall et al., 2018] and UCR [Dau et al., 2019] time series datasets, as well as some other time series classical datasets [Zhang et al., 2017, Wang et al., 2024b, Liu et al., 2022b, McCracken and Ng, 2016, Taieb et al., 2012]. The final list of all pre-training datasets is shown in Table 6. There is no overlap between the pre-training datasets and the target datasets. It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task. The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web. The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world. For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
@@ A.1.2 Evaluation Datasets
Table 6: List of pretraining datasets.
Domain Dataset Frequency Time Pionts Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash[Godahewa et al., 2021]
Wind 4 Seconds 7397147 Monash[Godahewa et al., 2021]
Wind Farms Minutely 172178060 Monash[Godahewa et al., 2021]
Solar 10 Minutes 7200720 Monash[Godahewa et al., 2021]
Solar Power 4 Seconds 7397222 Monash[Godahewa et al., 2021]
London Smart Meters Half Hourly 166527216 Monash[Godahewa et al., 2021]
Nature Phoneme - 2160640 UCRDau et al. [2019]
EigenWorms - 27947136 UEA[Bagnall et al., 2018]
PRSA Hourly 4628448 [Zhang et al., 2017]
Temperature Rain Daily 23252200 Monash[Godahewa et al., 2021]
StarLightCurves - 9457664 UCR[Dau et al., 2019]
Worms 0.033 Seconds 232200 UCR[Dau et al., 2019]
Saugeen River Flow Daily 23741 Monash[Godahewa et al., 2021]
Sunspot Daily 73924 Monash[Godahewa et al., 2021]
Weather Daily 43032000 Monash[Godahewa et al., 2021]
KDD Cup 2018 Daily 2942364 MonashGodahewa et al. [2021]
US Births Daily 7305 Monash[Godahewa et al., 2021]
Health MotorImagery 0.001 Seconds 72576000 UEA[Bagnall et al., 2018]
SelfRegulationSCP1 0.004 Seconds 3015936 UEA[Bagnall et al., 2018]
SelfRegulationSCP2 0.004 Seconds 3064320 UEA[Bagnall et al., 2018]
AtrialFibrillation 0.008 Seconds 38400 UEA[Bagnall et al., 2018]
PigArtPressure - 624000 UCR[Dau et al., 2019]
PIGCVP - 624000 UCR[Dau et al., 2019]
TDbrain 0.002 Seconds 79232703 [Wang et al., 2024b]
Transport Pems03 5 Minute 9382464 [Liu et al., 2022b]
Pems04 5 Minute 5216544 [Liu et al., 2022b]
Pems07 5 Minute 24921792 [Liu et al., 2022b]
Pems08 5 Minute 3035520 [Liu et al., 2022b]
Pems-bay 5 Minute 16937700 [Liu et al., 2022b]
Pedestrian_Counts Hourly 3132346 Monash[Godahewa et al., 2021]
Web Web Traffic Daily 116485589 Monash[Godahewa et al., 2021]
Economic FRED_MD Monthly 77896 [McCracken and Ng, 2016]
Bitcoin Daily 75364 Monash[Godahewa et al., 2021]
NN5 Daily 87801 [Taieb et al., 2012]
@@ A.1.3 Setting
We implemented ROSE in PyTorch [Paszke et al., 2019] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU. We used ADAM [Kingma and Ba, 2014] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training. By default, ROSE contains 3 encoder layers and 3 decoder layers with head number of 16 and the dimension of latent space $D=256$. The patch size for patching is set to 64.
@@ A.1.4 Baselines
We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer [Liu et al., 2023], PatchTST [Nie et al., 2022], TimesNet [Wu et al., 2022], and DLinear [Zeng et al., 2023], and two LLM-based models: GPT4TS [Zhou et al., 2024] and $\textbf{S}^{2}$-LLM [Pan et al., 2024]. In addition, we selected five foundation models for comparison in zero-shot setting, including Timer [Liu et al., 2024], MOIRAI [Woo et al., 2024], Chronos [Ansari et al., 2024], TimesFM [Das et al., 2023b] and Moment [Goswami et al., 2024]. The specific code base for these models is listed in Table 7:
@@
S2P-LLM https://github.com/panzijie825/S2IP-LLM
@@ A.6 Short-term forecasting
We also try to apply ROSE to short-term forecasting on the M4 [Makridakis, 2018] dataset, which contains the yearly, quarterly and monthly collected univariate marketing data. We follow TimesNet’s [Wu et al., 2022] setting and metrics (SMAPE, MASE and OWA) for testing. As shown in Table 9, ROSE also exhibits competitive performance on the M4 dataset compared to the baselines.
@@ A.7 Model generality
We evaluate the effectiveness of our proposed multi-frequency masking on Transformer-based models and CNN-based models, whose results are shown in Table 10. It is notable that multi-frequency masking consistently improves these forecasting models. Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [Wu et al., 2021], TimesNet [Wu et al., 2022], and PatchTST [Nie et al., 2022], respectively. This indicates that multi-frequency Masking can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
@@
Autoformer +Multi-frequency Masking 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
@@
TimesNet +Multi-frequency Masking 0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
@@
PatchTST +Multi-frequency Masking 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381

### v3

@@ 4 Experiments
Pre-training datasets. The datasets are crucial for pre-training a general time series forecasting model. In light of this, we gather many publicly available datasets from various domains, including energy, nature, health, transport, web, economics, etc. The details of these datasets are shown in the Appendix A.1.1. To enhance data utilization, we downsample fine-grained datasets to coarser granularity, resulting in approximately 887 million time points.
Evaluation datasets. To conduct comprehensive and fair comparisons for different models, we conduct experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
Baselines. We select the state-of-the-art models as baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$IP-LLM (Pan et al., 2024). In addition, we select five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b), and Moment (Goswami et al., 2024).
Setup. Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. Due to ROSE mostly aims at long-term predictions, for fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$. More implementation details are presented in the Appendix A.1.3.
@@ 4.1 In Distribution Forecasting
Setting. In full-shot setting, we utilize full downstream data to fine-tune pre-trained ROSE and baselines. In few-shot setting, we fine-tune all models with only 10% train data. The "Drop Last" issue is reported by several researchers (Qiu et al., 2024, 2025a; Li et al., 2025). That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size. In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
@@
Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
@@
Few-shot results. The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2. ROSE outperforms advanced models when training data is scarce in the target domain. Figure 4 shows the performance of pre-trained ROSE and ROSE trained from scratch on ETTh1 and ETTm2 with different fine-tuning data percentages, noting the best baselines in full-shot setting. The pre-trained ROSE shows stable, superior performance even with limited fine-tuning samples. Specifically, the pre-trained ROSE exceeds SOTA performance with only 1% train data for ETTh1 and 2% for ETTm2. Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in prediction performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
@@ 4.2 Zero-shot Forecasting
Results. As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in MSE. In comparison to Timer and Moirai, ROSE achieves average MSE reductions of 9% and 6%, respectively, and demonstrates a remarkable 43% relative improvement over Moment. Notably, ROSE stands out not only for its superior performance but also for its exceptionally lightweight and efficient design, which sets it apart from other foundational models. Detailed analysis of these aspects will be presented in Section 4.3.
Table 2: The results for ROSE and other foundation models in the zero-shot setting. The average results of all predicted lengths are listed here. We use ’-’ to indicate that the dataset has been involved in the model’s pre-training, and thus not used for testing.
@@
Metric MSE MSE MSE MSE MSE MSE
ETTh1 0.401 0.451 0.475 0.560 0.489 0.708
ETTh2 0.346 0.366 0.379 0.392 0.396 0.392
ETTm1 0.525 0.544 0.714 0.636 0.434 0.697
ETTm2 0.299 0.360 0.343 0.313 0.320 0.319
Weather 0.265 0.292 0.267 0.288 - 0.291
Electricity 0.234 0.297 0.241 0.245 - 0.861
Traffic 0.588 0.613 - 0.615 - 1.411
@@ 4.3 Model Analysis
As shown in Figure 5, ROSE is a lightweight general model with 7.4M parameters and short inference time, which are only about one-tenth of the second fastest/smallest foundation model (Timer). Importantly, ROSE uses the least number of parameters among foundation models, with its parameter count approaching that of specific models, while exhibiting superior zero-shot performance. This is attributed to our proposed decomposed frequency learning that enhances the comprehension of time series. Concurrently, the TS-Register achieves the adaptive transfer thus efficiently adapting to downstream tasks without the need of scaling up to achieve strong generalizability. Compared to foundation models with large scale, ROSE may better meet the need for general models in real scenarios that require high computational and parameter efficiency as well as high prediction accuracy with scarce downstream data.
@@ 4.4 Ablation Studies
Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
@@ 5 Conclusion and Future Work
In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance. ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios. Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities. Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification. We provide our code at https://github.com/decisionintelligence/ROSE.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
This work was partially supported by National Natural Science Foundation of China (62372179). Chenjuan Guo is the corresponding author of the work.
@@ A.1.1 Pre-training Datasets
We use multi-source datasets in pre-training which contain subsets of Monash (Godahewa et al., 2021), UEA (Bagnall et al., 2018) and UCR (Dau et al., 2019) time series datasets, as well as some other time series classical datasets (Zhang et al., 2017; Wang et al., 2024; Liu et al., 2022a; McCracken & Ng, 2016; Taieb et al., 2012). The final list of all pre-training datasets is shown in Table 5. There is no overlap between the pre-training datasets and the target datasets. It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task. The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web. The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world. For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
Table 5: List of pretraining datasets.
Domain Dataset Frequency Time Pionts Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash(Godahewa et al., 2021)
Wind 4 Seconds 7397147 Monash(Godahewa et al., 2021)
Wind Farms Minutely 172178060 Monash(Godahewa et al., 2021)
Solar 10 Minutes 7200720 Monash(Godahewa et al., 2021)
Solar Power 4 Seconds 7397222 Monash(Godahewa et al., 2021)
London Smart Meters Half Hourly 166527216 Monash(Godahewa et al., 2021)
Nature Phoneme - 2160640 UCR(Dau et al., 2019)
EigenWorms - 27947136 UEA(Bagnall et al., 2018)
PRSA Hourly 4628448 (Zhang et al., 2017)
Temperature Rain Daily 23252200 Monash(Godahewa et al., 2021)
StarLightCurves - 9457664 UCR(Dau et al., 2019)
Worms 0.033 Seconds 232200 UCR(Dau et al., 2019)
Saugeen River Flow Daily 23741 Monash(Godahewa et al., 2021)
Sunspot Daily 73924 Monash(Godahewa et al., 2021)
Weather Daily 43032000 Monash(Godahewa et al., 2021)
KDD Cup 2018 Daily 2942364 Monash(Godahewa et al., 2021)
US Births Daily 7305 Monash(Godahewa et al., 2021)
Health MotorImagery 0.001 Seconds 72576000 UEA(Bagnall et al., 2018)
SelfRegulationSCP1 0.004 Seconds 3015936 UEA(Bagnall et al., 2018)
SelfRegulationSCP2 0.004 Seconds 3064320 UEA(Bagnall et al., 2018)
AtrialFibrillation 0.008 Seconds 38400 UEA(Bagnall et al., 2018)
PigArtPressure - 624000 UCR(Dau et al., 2019)
PIGCVP - 624000 UCR(Dau et al., 2019)
TDbrain 0.002 Seconds 79232703 (Wang et al., 2024)
Transport Pems03 5 Minute 9382464 (Liu et al., 2022a)
Pems04 5 Minute 5216544 (Liu et al., 2022a)
Pems07 5 Minute 24921792 (Liu et al., 2022a)
Pems08 5 Minute 3035520 (Liu et al., 2022a)
Pems-bay 5 Minute 16937700 (Liu et al., 2022a)
Pedestrian_Counts Hourly 3132346 Monash(Godahewa et al., 2021)
Web Web Traffic Daily 116485589 Monash(Godahewa et al., 2021)
Economic FRED_MD Monthly 77896 (McCracken & Ng, 2016)
Bitcoin Daily 75364 Monash(Godahewa et al., 2021)
NN5 Daily 87801 (Taieb et al., 2012)
@@ A.1.3 Setting
We implemented ROSE in PyTorch (Paszke et al., 2019) and all the experiments were conducted on 8 NVIDIA A800 80GB GPU. We used ADAM (Kingma & Ba, 2014) with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training. By default, ROSE contains 3 encoder layers and 3 decoder layers with head number of 16 and the dimension of latent space $D=256$. The patch size for patching is set to 64.
@@
The t-SNE visualization. We select three datasets (Pems08, PSRA, Electricity) from transport, nature and energy domains respectively and compare the differences in hidden representations between direct transfer and adaptive transfer. Specifically, direct transfer refers to the case where domain specific information is not considered, while adaptive transfer considers domain specific information that is learned by register tokens. We visualized the output of the encoder’s hidden representations using t-SNE.
@@ A.1.4 Baselines
We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$IP-LLM (Pan et al., 2024). In addition, we selected five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b) and Moment (Goswami et al., 2024). The zero-shot experiment of Moment is designed based on the reconstruction task in the pre-training phase. Specifically, to ensure consistency between pre-training (reconstruction) and downstream prediction tasks, we actively mask the time periods to be predicted in the input sequence, and directly use the model’s reconstruction output for this part as the prediction value. Moment itself is not designed for zero-shot prediction and does not officially support zero-shot forecasting in this manner. The specific code base for these models is listed in Table 7:
@@
S2IP-LLM https://github.com/panzijie825/S2IP-LLM
@@ A.6 Short-term Forecasting
We also try to apply ROSE to short-term forecasting on the M4 (Makridakis, 2018) dataset, which contains the yearly, quarterly and monthly collected univariate marketing data. We follow TimesNet’s (Wu et al., 2022) setting and metrics (SMAPE, MASE and OWA) for testing. As shown in Table 9, ROSE also exhibits competitive performance on the M4 dataset compared to the baselines.
@@ A.7 Model Generality
We evaluate the effectiveness of our proposed multi-frequency masking on Transformer-based models and CNN-based models, whose results are shown in Table 10. It is notable that multi-frequency masking consistently improves these forecasting models. Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer (Wu et al., 2021a), TimesNet (Wu et al., 2022), and PatchTST (Nie et al., 2022), respectively. This indicates that multi-frequency Masking can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
@@
Autoformer
+Multi-frequency Masking
0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
@@
TimesNet
+Multi-frequency Masking
0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
@@
PatchTST
+Multi-frequency Masking
0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381